OpenStudy (anonymous):

Fans & Medals!! Solve the equation 5x – (3x – 2) = 28 Solve the equation 5x – (3x – 2) = 28 x = 13 x = 14 x = 15 x = 16

OpenStudy (anonymous):

first add (3x-2) to both sides. it cancels out on the left and you now have \[5x=28+3x+2\] Now combine like terms. Tell me what you have after you combine like terms.

OpenStudy (anonymous):

Hint: add the 28 and the 2 together.

OpenStudy (kendricklamar2014):

5x−(3x−2)=28 Step 1: Simplify both sides of the equation. 2x+2=28 Step 2: Subtract 2 from both sides. 2x+2−2=28−2 2x=26 Step 3: Divide both sides by 2. 2x 2 = 26 2 x=13

OpenStudy (anonymous):

@KendrickLamar2014 http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

"Give Help, Not Answers - I will encourage and guide those needing help, and not just give them an answer"

OpenStudy (anonymous):

Thank you @KendrickLamar2014

OpenStudy (anonymous):

He's wrong.

OpenStudy (anonymous):

\[5x=30+3x\] Subtract 3x from both sides. \[2x=30\] divide both sides by 2 \[x=15\]